OpenStudy (anonymous):

Phillip bought 12 used CDs and DVDs. CDs cost $2 each and DVDs cost $3 each. He spent $31 not including tax. How many DVDs did Phillip buy?

OpenStudy (anonymous):

alright you know so far this much, x(cd)+y(dvd)=12(total bought) and 2x+3y=31, so what you do first is solve for the first portion, for example subtract y from x=12 so it will become x=12-y, than you substitute 2(x) with 2(12-y)+3y=31. Distribute that 2 and I think you should be able to take it from there.